a train travels 50 km in half hour. it the stops at a station for 20 mins, before traveling for 2 hours at an average speed of 90 km/h what was the trains average speed over the whole journey
1
Expert's answer
2020-08-14T16:01:35-0400

By definition, the average speed is:


"v_{av} = \dfrac{l_{tot}}{T_{tot}}"

where "l_{tot}" is the total distance covered by the train in time "T_{tot}".

The distance covered during the first part of the journey is 50 km, during the second part is 0 km and during the third part "90km/h\cdot 2h = 180" km. Thus:


"l_{tot} = 50 km + 0km + 180km = 230km"

The time spent on the first part of the journey is 0.5h, on the second part "20/60 = 1/3 \space h" and on the third part 2h, The total time is:


"T_{tot} = 0.5h + 1/3\space h + 2h = 17/6 \space h"

The average speed wil be:


"v_{av} = \dfrac{l_{tot}}{T_{tot}} = \dfrac{230km}{17/6\space h} \approx 81.2 \space km/h"

Answer. 81.2 km/h.